1. The boiling point of an aqueous solution containing a non-volatile electrolyte is 100.75 degrees C. 
  2. What is the boiling point elevation? 
  3. What is the molality of the solution?
1
Expert's answer
2021-05-11T05:55:49-0400

Solution:

A solution will boil at a higher temperature than the pure solvent. This is the colligative property called boiling point elevation.

An equation has been developed for this behavior. It is:

Δt = i × Kb × m

Δt = the boiling point elevation

Kb for water is 0.512°C kg mol¯1

i = 1 (for non-volatile electrolyte)

m = the molality of the solute


The boiling point of pure water is 100°C.

Thus,

Δt = 100.75°C - 100.00°C = 0.75°C

Δt = 0.75°C


m = Δt / (i × Kb) = (0.75°C) / (1 × 0.512°C kg mol¯1) = 1.4648 mol/kg = 1.465 mol/kg

m = 1.465 mol/kg


Answers:

The boiling point elevation (Δt) is 0.75°C

The molality of the solution (m) is 1.465 mol/kg